Evaluate the iterated integral ∫(0 to 1)∫(x to 5x)∫(0 to y) 5xyz dzdydx.
1
Expert's answer
2020-11-09T15:32:41-0500

"\int_0^1 \int_x^{5x} \int_0^y 5xyz dzdydx"


Go from inside to outside


"1. \int_{0}^{y} 5xyz dz = 5xy\Big[z^2/2\Big]_0^y = 5/2xy^3"


"2. \int_{x}^{5x} 5/2xy^3 dy = 5/2 x\Big[y^4/4\Big]_x^{5x} = 5/2x(625x^4/4 - x^4/4) = 390x^5"


"3. \int_0^1 390x^5 dx = 390x^6/6\Big|_0^1 = 65"


Answer: "\int_0^1 \int_x^{5x} \int_0^y 5xyz dzdydx = 65"